OpenStudy (anonymous):

The height of an object thrown vertically upward from a height of 25 feet is given by h(t) = -16t^2 + 128t +25 where the height is measured in feet and time is measured in seconds. How long does it take for the object to hit the ground?

OpenStudy (anonymous):

you have a curious mind...

OpenStudy (anonymous):

Physics

OpenStudy (anonymous):

@vikayefremova

OpenStudy (anonymous):

@ParthKohli @lgbasallote @lalaly

OpenStudy (anonymous):

my job is over

OpenStudy (anonymous):

over to you guys

Parth (parthkohli):

So, we know that the height must be 0(it's the ground right?) Solve this quadratic equation. \( \color{Black}{\Rightarrow -16t^2 + 128 + 25 = 0}\) Do you know how to solve quadratic equations?

OpenStudy (anonymous):

I forgot, we learned this so long ago. o;

Parth (parthkohli):

Use the quadratic formula here. \( \color{Black}{\Rightarrow t = \Large {-b \pm\sqrt{b^2 - 4ac} \over 2a} }\)

Parth (parthkohli):

a = -16 b = 128 c = 25 Use a calculator to do this.

OpenStudy (anonymous):

Am I supposed to be getting huge numbers

Parth (parthkohli):

No...not at all

OpenStudy (anonymous):

Show me what you did please?

OpenStudy (anonymous):

Ok assume you throw a ball upwards, it will get to a maximum height and then it will come down again right? Reply Using Drawing This is how it's path would look like Forming a parabola. So when the ball comes back down, and hits the ground the height of the ball will be zero since it's on the ground. So to find the time it takes to hit the ground, we find the time when the height is equal to zero. So you just put h as o and solve for t. Factoring out -16t and then you would have -16t(t-4) = 0 Equations of such form give you two answers, 16t=0 which is just t=0 and t-4=0 which is t=4 I hope that helps. This is a similar question

OpenStudy (anonymous):

But for yours, using the quadratic formula, you can solve for t, Just remember a= coefficient of x^2 b= coefficient of x and c = constant

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay, thanks. o; But like, maybe I'm setting up the quadratic formula wrong which I shouldn't be.

OpenStudy (anonymous):

t= \[(-128+/- \sqrt{(128)^{2}-4(-16\times25}))/2\times-16\] That's what you should have

OpenStudy (anonymous):

\[-128 \pm \sqrt{17225}/-32\]

OpenStudy (anonymous):

Couldn't that be simplified more? o:

OpenStudy (anonymous):

And wouldn't that add that imaginary thing..? idk

OpenStudy (anonymous):

\[4±\sqrt{281}/-4\] This is how it would be, ignore my previous one

OpenStudy (anonymous):

She told us that the answer is 8.19 seconds, so how would we get from that to there?

OpenStudy (anonymous):

But then remember, time can not be negative, so you're answer would be \[4-\sqrt{281}/-4\] Notice that if you try to do \[4+\sqrt{281}/-4\] it gives you a negative answer?

OpenStudy (anonymous):

Mhm. But I still don't know how to get 8.19 seconds. :c

OpenStudy (anonymous):

ok \[\sqrt{281}\] using a calculator is around 16.7631 if you divide that by -4, you get -4.1908. So now all you need to do is 4-(-4.1908)= 4+4.1908=8.19

OpenStudy (anonymous):

Ohohoh. Okay. C: Tysm.